silvernight269:

find the constant difference for each of the tables below. use these results to determine if each table represents a linear, quadratic, or a cubic function. then find the equation of each function.

silvernight269:

1 attachment
silvernight269:

@dude

dude:

Sorry for the late response First find the slope for the first table Chose any 2 points (preferably ones next to each other) Say for example you choose (-13,41) and (-7,23) ( x\(_1\),\(y_1\)) (x\(_2\), \(y_2\)) Use the equation to find slope \(\large\frac{y_2-y_1}{x_2-x_1}=\frac{23-41}{-7-(-13)}\) Once you do that choose another two points Lets say you choose (-7,23) and (-1,5) ( x\(_1\),\(y_1\)) (x\(_2\), \(y_2\)) And use the same equation \(\large\frac{y_2-y_1}{x_2-x_1}=\frac{5-23}{-1-(-7)}\) If they are the same, then it is a linear function (they have the same slope)

dude:

From them you can just use any point (from the table) and substitute what you have to find the y-intercept

dude:

For the second table we can clearly see that they are not going on a constant rate In terms of it being a quadratic function, the points would be the same on both positive and negative values after it reaches its vertex

dude:

So it is automatically a cubic function (I am not sure how to fully explain this however here is the graph and so I hope it does help, will try to explain though) https://www.desmos.com/calculator/uu5ygqwf24